# Hierarchical Multi-agent Reinforcement Learning for Cyber Network Defense

### v1

@@ 3. Problem Statement
In cyber security, red teams act as attackers who attempt to exploit network vulnerabilities and carry out malicious activities aimed at compromising the system. Blue teams are tasked with defending against red team opponents to secure the networks, while maintaining network operations. In this work, we focus on the CybORG CAGE 4 cybersecurity MARL framework Group (2023), which is a realistic environment that models cyber defense. We discuss several aspects of the environment below. For additional details, please see Section 1 in the supplemental materials, and the CAGE 4 description Group (2023).
@@ 4.1. Hierarchical MARL Design
In the CAGE 4 setup, blocking the traffic between neighboring subnets is not particularly useful for two reasons: (i) it prevents green agents from accessing remote services, which penalizes the blue team, and (ii) red agents are active mostly within their own assigned subnet, and rarely reach remotely into other subnets, a setup that ensures scaling with large networks. In the rest of the paper, unless otherwise specified, we discuss defenses that do not use block or allow traffic actions, focusing on the Investigate and Recover sub-policies. We explore traffic control guided by communication among blue agents in the supplemental material.
@@ 5.3. H-MARL Transferability
Figure 6 presents the fine-tuning results of the Investigate sub-policies, pre-trained against Aggressive Red, the average-performing attacker. A short fine-tuning is enough to adapt to a new adversary, resulting in learning curves similar to those obtained when training from scratch. The fine-tuning results for the Recover sub-policy are included in Section 2 of the supplemental material. We note that Investigate learns different strategies for each adversary, while Recover is relatively agnostic to the attack, and can be directly reused against other red agents. This is understandable, given that the frequency of analyze actions or the placement of decoys is affected by how stealthy or impactful an attacker is, while the recover strategy is attack agnostic, always targeting hosts that present indicators of compromise and performing a remove or restore action.
@@ 5.5. What method to use: Expert or Meta?
In our previous experiments, we have seen that H-MARL Expert performs best, as it is guided by a deterministic master policy generated from domain knowledge, specifically: If host IOCs are present, invoke the Recover sub-policy; otherwise Investigate. To show that such expert rules do not generalize in all situations, we consider a scenario with a third Control Traffic sub-policy, in which a defender may block traffic between a clean subnet and another subnet that presents IOCs. However, block actions incur penalties for preventing the remote activity of green users, rendering a deterministic blocking rule unfavorable. We experimented with a probabilistic master policy instead: If host IOCs are present, invoke the Recover sub-policy; otherwise split 75%-25% between Investigate and Control Traffic. With this probabilistic rule, the H-MARL Expert method becomes unstable, while H-MARL Meta performs better, converging fast to a stable performance (Figure 9). These experiments demonstrate the importance of learning the master policy and are discussed in more detail in the supplemental material, Section 4.
@@ 6. Conclusion
Appendix A Communication
@@ Appendix B Traffic Control
Turning off Degrade Service. In our next set of experiments illustrated in Figure 10, we fine-tuned the sub-policies after turning off the reward penalty of green agents being affected by failed service access. We also used the modified red agent introduced in Section A of the supplemental material, which performs remote scanning into other subnets more often, and can still collect rewards by impacting the critical OT service. With block actions being now useful at preventing red agents from spreading, without incurring penalties, H-MARL Expert regains a stable, high performance, as expected. H-MARL Meta achieves similar performance with H-MARL Expert in this setting. The blue agents are using 4×4$\times$4 × more block actions when the failed user access penalty has been removed.

### v2

@@ 3. Problem Statement
In cyber security, red teams act as attackers who attempt to exploit network vulnerabilities and carry out malicious activities aimed at compromising the system. Blue teams are tasked with defending against red team opponents to secure the networks, while maintaining network operations. In this work, we focus on the CybORG CAGE 4 cybersecurity MARL framework Group (2023), which is a realistic environment that models cyber defense. We discuss several aspects of the environment below. For additional details, please see Section A.1 in the supplemental materials, and the CAGE 4 description Group (2023).
@@ 4.1. Hierarchical MARL Design
In the CAGE 4 setup, blocking the traffic between neighboring subnets is not particularly useful for two reasons: (i) it prevents green agents from accessing remote services, which penalizes the blue team, and (ii) red agents are active mostly within their own assigned subnet, and rarely reach remotely into other subnets, a setup that ensures scaling with large networks. In the rest of the paper, unless otherwise specified, we discuss defenses that do not use block or allow traffic actions, focusing on the Investigate and Recover sub-policies. We explore traffic control guided by communication among blue agents in supplemental material Section A.4.
@@ 5.3. H-MARL Transferability
Figure 6 presents the fine-tuning results of the Investigate sub-policies, pre-trained against Aggressive Red, the average-performing attacker. A short fine-tuning is enough to adapt to a new adversary, resulting in learning curves similar to those obtained when training from scratch. The fine-tuning results for the Recover sub-policy are included in Section A.2 of the supplemental material. We note that Investigate learns different strategies for each adversary, while Recover is relatively agnostic to the attack, and can be directly reused against other red agents. This is understandable, given that the frequency of analyze actions or the placement of decoys is affected by how stealthy or impactful an attacker is, while the recover strategy is attack agnostic, always targeting hosts that present indicators of compromise and performing a remove or restore action.
@@ 5.5. What method to use: Expert or Meta?
In our previous experiments, we have seen that H-MARL Expert performs best, as it is guided by a deterministic master policy generated from domain knowledge, specifically: If host IOCs are present, invoke the Recover sub-policy; otherwise Investigate. To show that such expert rules do not generalize in all situations, we consider a scenario with a third Control Traffic sub-policy, in which a defender may block traffic between a clean subnet and another subnet that presents IOCs. However, block actions incur penalties for preventing the remote activity of green users, rendering a deterministic blocking rule unfavorable. We experimented with a probabilistic master policy instead: If host IOCs are present, invoke the Recover sub-policy; otherwise split 75%-25% between Investigate and Control Traffic. With this probabilistic rule, the H-MARL Expert method becomes unstable, while H-MARL Meta performs better, converging fast to a stable performance (Figure 10). These experiments demonstrate the importance of learning the master policy and are discussed in more detail in the supplemental material, Section A.4.
@@ 6. Conclusion
Appendix A Supplemental Material
Here we provide some additional experiments and context supplemental to the main body of the paper. In A.1 we provide further details on the CAGE-4 cyber game studied in this paper, in A.2 we show additional transferability experiments on H-MARL Meta, in A.3 we study the potential uses of communication between agents, and in A.4 we further explore the use of Block and Allow Traffic actions in CAGE-4.
A.1. The cyber game
This section presents additional details on the cyber game simulated in CybORG CAGE 4  [^1] https://github.com/cage-challenge/cage-challenge-4/tree/main.
For increased realism, users on enterprise machines are represented by green agents, which are present on every host. Green agents randomly access local and remote services, such as SSHD, MySQL, FTP, etc. Upon compromise, red agents are able to degrade these services, preventing users from completing their work or slowing them down.
The red team can attack the network through several actions: scan the network to discover new hosts; scan a host to discover active services; exploit a vulnerability to compromise a host; escalate privileges on a compromised node to gain root access; degrade user experience of green agents; impact (stop) the critical Operational technology (OT) service; discover deception (i.e., probe a host to determine if it is running decoy services).
The action sets for green agents (users), blue agents (defenders) and red agents (attackers) are presented in Table 4. Both blue and red agent actions have a specified time duration, ranging from one time step (block/allow traffic) to five time steps for longer actions (restore host). Agents are not allowed to choose another action before the previous one has completed.
Figure 8. The Recover sub-task is rather agnostic to the attack type and can be re-used against other adversaries.
Table 4. Action sets for green agents (users), blue agents (defenders), and red agents (attackers).
Green actions
Green Access Service Communicate with a server in the local zone or another zone. It has a small chance of being flagged as malicious by Monitor.
Green Local Work Do work on the local host without communicating. It has a small chance that this action results in red gaining a foothold on the host (e.g. as a result of a Phishing email). Also, it has a small chance of being flagged as malicious by Monitor.
Blue actions
Monitor Collection of information about flagged malicious activity on the network.
Analyse Collection of further information relating to malware files, on a specific host.
Deploy Decoy Setup of a decoy service (type randomly assigned) on a specified host. Blue is alerted if Red accesses this service.
Remove Attempts to stop all processes identified as malicious by the monitor action.
Restore Restoring a system to a known good state.
Block Traffic Block traffic to and from the specified zone.
Allow Traffic Allow traffic to and from the specified zone.
Red actions
Discover Remote Systems Discovers new hosts/IP addresses in the network through active scanning using tools such as ping.
Aggressive Service Discovery Discovers responsive services on a selected host by initiating a connection with that host.
Stealth Service Discovery Same as Aggressive Service Discovery but slower and quieter (i.e., lower chance of raising alerts).
Exploit Network Services Attempts to exploit a specified service on a remote system.
Privilege Escalate This action escalates the agent’s privilege on the host.
Impact This action disrupts the performance of the network and fulfils red’s objective of denying the operational service.
Degrade Services If red has root privileges on a host, it may degrade the user experience for a green agent.
Discover Deception Probe a host to determine if it is running decoy services.
Withdraw Remove red presence from target host.
A.2. H-MARL Transferability
Figure 8 presents the fine-tuning results of the Recover sub-policy, pre-trained against Aggressive Red, the average-performing attacker. The Recover sub-policy is trained on an observation space consisting of indicators of compromise within a subnet. This sub-policy learns a strong strategy regardless of the attack, and can be directly reused against other red agents.
A.3. Communication
@@ A.4. Traffic Control
Turning off Degrade Service. In our next set of experiments illustrated in Figure 11, we fine-tuned the sub-policies after turning off the reward penalty of green agents being affected by failed service access. We also used the modified red agent introduced in Section A.3 of the supplemental material, which performs remote scanning into other subnets more often, and can still collect rewards by impacting the critical OT service. With block actions being now useful at preventing red agents from spreading, without incurring penalties, H-MARL Expert regains a stable, high performance, as expected. H-MARL Meta achieves similar performance with H-MARL Expert in this setting. The blue agents are using 4×4$\times$4 × more block actions when the failed user access penalty has been removed.